OpenStudy (anonymous):

Select the best possible first step to solving the following system by first eliminating the y variable. 3x + y = 6 x + 2y = –13 Multiply the first equation by 2. Multiply the second equation by 2. Multiply the first equation by –2. None of the above.

OpenStudy (anonymous):

solve for y in the first equation. subtract 3x from both sides

OpenStudy (anonymous):

Multiply the first eq by 2. so you will get \[6x + 2y = 12, x + 2y = -13\]

OpenStudy (anonymous):

then subtract the second from the first, you will get : \[5x = 25, \therefore x = 5. \]

OpenStudy (anonymous):

then put x in the ORIGINAL eq, so you will get: \[2y = -18, \therefore y = -9\]

OpenStudy (anonymous):

wrong it wants to eliminate y first not x. solve for y in the first then use that to substitute into the second equation.

OpenStudy (anonymous):

dudee. i eliminated y. you can't see y when it's gone, or as aptly put, eliminated.

OpenStudy (anonymous):

y=6-3x, x+2(6-3x)=-13. the question says eliminate y first. you eliminated x first.

OpenStudy (anonymous):

your final answer was correct but you failed to read the question correctly.

OpenStudy (anonymous):

nononononon. see this. you are using the subsitution method, he asked for the elimination method:  see that there??? that, is y being taken out, or eliminated.